# Kinetic study of Kelvin-Helmholtz instability: Multiscale thermodynamic nonequilibrium effects and their relative importance

### v1

@@ IV.2.3 Comparison of Macroscopic Quantities
Figure 14 shows the distributions of macroscopic quantities computed by both models. Macroscopic quantities play a fundamental role in traditional fluid mechanics. During the evolution of the KHI, the relative error in macroscopic quantities computed by the low-order model accumulates over time. As second-order nonequilibrium effects become significant, a complex interplay arises, with first- and second-order nonequilibrium effects alternating. This results in substantial errors in the macroscopic quantities computed by the D2V16 model. At $t=1.68$, when large vortices fully develop, the accumulated errors in macroscopic quantities computed by D2V16 reach their peak. The relative errors in density, temperature, and pressure reach $43\%$, while the velocity error exceeds $200\%$ at $x\approx 180$, as shown in Fig. 14(d).
@@
(2) Errors in higher-order moment relations: The primary limitation of D2V16 is its failure to preserve higher-order moment relations. Errors in higher-order moments propagate through constitutive relations, impacting lower-order moments and, consequently, macroscopic quantities.
These findings further confirm that, in complex fluid systems with significant second-order nonequilibrium effects, low-order models fail to provide adequate descriptions. Therefore, high-order models are essential for accurate simulations.
@@ IV.3 Dominance of second-order TNEs
$\bm{\Xi}_{\text{TNE}}=\frac{\tau\Delta\mathbf{u}}{\Delta\mathbf{r}}.$
@@
IV.4 Effectiveness, extensibility, and limitations of high-order models
The next step is to further increase ${\tau}/{\Delta x}$ and evaluate the upper limit of the D2V25 model’s ability to describe nonequilibrium effects when second-order nonequilibrium effects fully dominate. The initial conditions are set as $\rho_{L}=6.0$, $\rho_{R}=2.0$, and $P_{L}=P_{R}=4.0$. The velocity perturbation in the $x$-direction is prescribed by Eq. (27), with an amplitude of $u_{0}=100\Delta y$ and an initial wavenumber $k=2\pi/L_{y}$. All other parameters, boundary conditions, and finite difference schemes remain the same as those used in the previous cases.
@@ IV.4.2 Effectiveness, extensibility, and limitations of D2V25 DVS
The D2V25 model in this study demonstrates excellent stability and accuracy in computing viscous stress during the evolution of the KHI. Next, we evaluate its ability to capture other nonequilibrium effects.
@@
Figure 21 shows the distributions of $\Delta_{3xxx}^{*}$ and $\Delta_{4,2xy}^{*}$ along $x=0.7L_{x}$ at $t=0.06$ and $t=1.2$. The nonequilibrium measures $\bm{\Delta}_{3}^{*}$ and $\bm{\Delta}_{4,2}^{*}$, along with their components, influence the accuracy of the constitutive relation evolution equations Gan et al. (2022), necessitating higher-order moment relations for precise computation. However, as shown in Fig. 21, the D2V25 DVS accurately captures higher-order nonequilibrium effects, including $\Delta_{3xxx}^{*}$ and $\Delta_{4,2xy}^{*}$, with numerical results closely aligning with the second-order analytical solutions. This demonstrates the capability of the D2V25 model to capture higher-order nonequilibrium effects, suggesting that its practical simulation performance exceeds its theoretical design to some extent.
@@
As shown in Fig. 22, the simulation results for $\Delta_{4yyyy}^{*}$ and $\Delta_{5,3xyy}^{*}$ using the D2V25 DVS are unsatisfactory. This indicates that while the D2V25 DVS demonstrates a certain level of extensibility, it remains inherently limited. To address higher-order nonequilibrium effects, additional moment relations for $f_{i}^{eq}$ are required, along with the adoption of higher-order models.
@@ V Conclusions, discussions, and outlooks
When second-order nonequilibrium effects come to the forefront, the battle between low- and high-order models becomes evident. Our comparative analysis reveals critical shortcomings in low-order models: they fall to accurately predict KHI growth rates, capture macroscopic evolution, preserve kinetic moment relationships, and faithfully reproduce thermodynamic nonequilibrium characteristics. In contrast, high-order models emerge as the superior contenders, offering sharper physical accuracy, greater cross-scale adaptability, and enhanced numerical stability in the presence of strong nonequilibrium effects. Notably, although high-order models exhibit some degree of extensibility in capturing strong nonequilibrium effects, their applicability remains fundamentally constrained. In physical modeling, the key lies in dynamically adjusting the preserved moment relations during phase-space discretization and tailoring discretization strategies to specific research objectives-only then can we ensure a precise and robust kinetic description. To further unravel the multiscale nature of KHI, we construct a nonequilibrium phase diagram-a roadmap that visually charts the interplay between different nonequilibrium orders. This diagram not only deepens our understanding of the instability’s intricate dynamics but also serves as a valuable guide for selecting suitable kinetic models across a broad spectrum of nonequilibrium conditions in both scientific research and engineering applications.
Despite these advances, several frontiers remain unexplored. First, our current model captures only second-order TNEs, whereas extreme conditions may demand the inclusion of third- and higher-order effects. Second, our analysis primarily focused on the role of viscous stress in shaping KHI. However, other nonequilibrium effects–driven by different physical mechanisms–may follow entirely different dominance pathways, which remain to be fully understood. Third, our study operates within the framework of the ideal gas assumption, whereas real-world scenarios introduce additional complexities such as surface tension and non-ideal gas effects, fundamentally altering the evolution of the instability.
To further push the boundaries of kinetic modeling further, future research will focus on developing a multiphase flow DBM that incorporates third- and higher-order TNEs. This model will faciliate the exploration of competition and coordination mechanisms among different orders and types of TNEs, driven by various forces such as surface tension and equation-of-state modifications during the development of KHI. This endeavor will not only advance the theoretical foundations of kinetic modeling but also pave the way for addressing real-world challenges in multiphase flows and nonequilibrium fluid dynamics.
@@ Acknowledgements
The authors sincerely thank Yanhong Wu, Zhaowen Zhuang and Jiahui Song for their valuable discussions and insights. This work was supported by the National Natural Science Foundation of China (Grant Nos. U2242214, 11875001, and 12172061), the Hebei Outstanding Youth Science Foundation (Grant No. A2023409003), the Central Guidance on Local Science and Technology Development Fund of Hebei Province (Grant No. 226Z7601G), the Fujian Provincial Units Special Funds for Education and Research (Grant No. 2022639), and the Foundation of the National Key Laboratory of Shock Wave and Detonation Physics (Grant No. JCKYS2023212003).

### v2

@@ IV.2.4 Comparison of macroscopic quantities
Figure 14 shows the distributions of macroscopic quantities computed by both models. Macroscopic quantities play a fundamental role in traditional fluid mechanics. During the evolution of the KHI, the relative error in macroscopic quantities computed by the low-order model accumulates over time. As second-order nonequilibrium effects become significant, a complex interplay arises, with first- and second-order nonequilibrium effects alternating. This results in substantial errors in the macroscopic quantities computed by the D2V16 DVS. At $t=1.68$, when large vortices fully develop, the accumulated errors in macroscopic quantities computed by D2V16 reach their peak. The relative errors in density, temperature, and pressure reach $43\%$, while the velocity error exceeds $200\%$ at $x\approx 180$, as shown in Fig. 14(d).
@@
(2) Errors in higher-order kinetic moments: The primary limitation of D2V16 is its failure to preserve higher-order kinetic moments. Higher-order moments refine physical models by capturing nonequilibrium effects beyond the NS framework. They accurately capture fine-scale fluid structures and rapid dynamics, directly influencing the characterization of nonequilibrium effects. The long-term accumulation of these effects significantly impacts macroscopic behavior.
The computational costs of high-order and low-order models are compared. For the same computational task, the ratio of computational time between the two models is 0.61, which closely matches the ratio of their discrete velocity counts (0.64). This indicates that computational efficiency is approximately inversely proportional to the number of discrete velocities. It is necessary to note that computational efficiency is not the primary criterion for model selection; rather, the corresponding physical functionality should guide the choice.
@@ IV.3 Dominance of second-order TNEs
$\bm{\Xi}_{\text{TNE}}=\frac{\tau\Delta\mathbf{u}}{\Delta\mathbf{r}}.$ (38)
@@
Compared with the traditional $Kn$, the criterion proposed in this study is based on $R_{\text{TNE}}$. Although both quantify the degree of nonequilibrium within a system, they differ in terms of descriptive perspective, sensitivity, and information richness. $Kn$ primarily relies on slow variables, such as macroscopic gradients, which makes it challenging to dynamically and precisely capture instantaneous changes and microscopic details within the fluid Gan et al. (2022). In contrast, $R_{\text{TNE}}$ reveals the relative significance of nonequilibrium effects at different orders through the ratios of nonequilibrium quantities, thereby offering a more accurate representation of the intrinsic characteristics of nonequilibrium states.
IV.4 Effectiveness, extensibility, and limitations of high-order DVS
The next step is to further increase ${\tau}/{\Delta x}$ and evaluate the upper limit of the D2V25 DVS’s ability to describe nonequilibrium effects when second-order nonequilibrium effects fully dominate. The initial conditions are set as $\rho_{L}=6.0$, $\rho_{R}=2.0$, and $P_{L}=P_{R}=4.0$. The velocity perturbation in the $x$-direction is prescribed by Eq. (27), with an amplitude of $u_{0}=100\Delta y$ and an initial wavenumber $k=2\pi/L_{y}$. All other parameters, boundary conditions, and finite difference schemes remain the same as those used in the previous cases.
@@ IV.4.2 Effectiveness, extensibility, and limitations of D2V25 DVS
The D2V25 DVS in this study demonstrates excellent stability and accuracy in computing viscous stress during the evolution of the KHI. Next, we evaluate its ability to capture other nonequilibrium effects.
@@
Figure 21 shows the distributions of $\Delta_{3xxx}^{*}$ and $\Delta_{4,2xy}^{*}$ along $x=0.7L_{x}$ at $t=0.06$ and $t=1.2$. The nonequilibrium measures $\bm{\Delta}_{3}^{*}$ and $\bm{\Delta}_{4,2}^{*}$, along with their components, influence the accuracy of the constitutive relation evolution equations Gan et al. (2022), necessitating higher-order kinetic moment relations for precise computation. However, as shown in Fig. 21, the D2V25 DVS accurately captures higher-order nonequilibrium effects, including $\Delta_{3xxx}^{*}$ and $\Delta_{4,2xy}^{*}$, with numerical results closely aligning with the second-order analytical solutions. This demonstrates the capability of the D2V25 DVS to capture higher-order nonequilibrium effects, suggesting that its practical simulation performance exceeds its theoretical design to some extent.
@@
As shown in Fig. 22, the simulation results for $\Delta_{4yyyy}^{*}$ and $\Delta_{5,3xyy}^{*}$ using the D2V25 DVS are unsatisfactory. This indicates that while the D2V25 DVS demonstrates a certain level of extensibility, it remains inherently limited. To address higher-order nonequilibrium effects, additional kinetic moment relations for $f_{i}^{eq}$ are required, along with the adoption of higher-order models.
@@ V Conclusions, discussions, and outlooks
When second-order nonequilibrium effects come to the forefront, the battle between low- and high-order models becomes evident. Our comparative analysis reveals critical shortcomings in low-order models: they fall to accurately predict KHI growth rates, capture macroscopic evolution, preserve kinetic moment relationships, and faithfully reproduce thermodynamic nonequilibrium characteristics. In contrast, high-order models emerge as the superior contenders, offering sharper physical accuracy, greater cross-scale adaptability, and enhanced numerical stability in the presence of strong nonequilibrium effects. Notably, although high-order models exhibit some degree of extensibility in capturing strong nonequilibrium effects, their applicability remains fundamentally constrained. In physical modeling, the key lies in dynamically adjusting the preserved kinetic moment relations during phase-space discretization and tailoring discretization strategies to specific research objectives-only then can we ensure a precise and robust kinetic description. To further unravel the multiscale nature of KHI, we construct a nonequilibrium phase diagram-a roadmap that visually charts the interplay between different nonequilibrium orders. This diagram not only deepens our understanding of the instability’s intricate dynamics but also serves as a valuable guide for selecting suitable kinetic models across a broad spectrum of nonequilibrium conditions in both scientific research and engineering applications.
Despite these advances, several frontiers remain unexplored. (i) Our current model captures only second-order TNEs, whereas extreme conditions may require the inclusion of third- and higher-order effects. (ii) Our analysis primarily focused on the role of viscous stress in shaping KHI. However, other nonequilibrium effects—driven by different physical mechanisms—may follow entirely different dominance pathways. For example, the evolution of heat flux follows distinct mechanisms: first-order heat flux is primarily driven by the temperature gradient, second-order heat flux involves the combined effects of temperature and velocity gradients, and third-order or higher-order heat flux is associated with more complex nonlinear effects. (iii) Our study is limited to the two-dimensional case. Expanding from two to three dimensions presents several challenges, including the design of a three-dimensional discrete velocity set, the theoretical derivation of nonequilibrium effects, and the analysis of three-dimensional physical fields. This paper does not analyze the impact of external fields on KHI. For instance, in astrophysics and plasma physics, electromagnetic field effects are often significant.
To further advance the boundaries of kinetic modeling, future research will focus on developing a three-dimensional DBM model that incorporates third-order and higher-order TNEs, as well as multiphysics coupling. This model will aid in exploring the competition and coordination mechanisms among different orders and types of TNEs during the evolution of KHI. Additionally, by modifying the external force term, electromagnetic field effects will be introduced to further investigate their influence on the development of KHI. These efforts will not only enrich the theoretical foundation of kinetic modeling and provide precise simulations and analyses in broader physical contexts, but also offer new insights and techniques for addressing complex phenomena in plasma physics, astrophysics, and related fields.
@@ Acknowledgements
The authors sincerely thank Yanhong Wu, Zhaowen Zhuang and Jiahui Song for their valuable discussions and insights. This work was supported by the National Natural Science Foundation of China (Grant Nos. U2242214, 52278119, 11875001, and 12172061), the Hebei Outstanding Youth Science Foundation (Grant No. A2023409003), the Central Guidance on Local Science and Technology Development Fund of Hebei Province (Grant No. 226Z7601G), the Fujian Provincial Units Special Funds for Education and Research (Grant No. 2022639), and the Foundation of the National Key Laboratory of Shock Wave and Detonation Physics (Grant No. JCKYS2023212003).